OpenStudy (anonymous):

Find the value of y when x equals 14 in this equation. -6x + 6y = -42

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

-6(14) + 6y = -42 -84 + 6y = -42 6y = 42 y = 7

OpenStudy (anonymous):

are you kiddin me? yu said 21

OpenStudy (anonymous):

i forgot that 42 was a negative, sorry.

OpenStudy (anonymous):

i'm takin a test && i got it wrong

OpenStudy (anonymous):

so -42 + 84 = 42

OpenStudy (anonymous):

too late now

OpenStudy (anonymous):

I would say, um... I would divide out the entire problem by 6 just to make it simpler. I am pretty sure its y = 7 as bunny stated though lol

OpenStudy (anonymous):

thnks